### v2

# Implicit Bias-Like Patterns in Reasoning Models

Abstract
Implicit bias refers to automatic mental processes that shape perceptions, judgments, and behaviors. Previous research on “implicit bias" in LLMs focused primarily on outputs rather than the processes underlying the outputs. We present the Reasoning Model Implicit Association Test (RM-IAT) to study implicit bias-like processing in reasoning models, which are LLMs using step-by-step reasoning for complex tasks. Using RM-IAT, we find o3-mini and DeepSeek R1 require more tokens when processing association-incompatible information, mirroring human implicit bias patterns. Conversely, Claude 3.7 Sonnet displays reversed patterns for race and gender tests, requiring more tokens for association-compatible information. This reversal appears linked to differences in safety mechanism activation, increasing deliberation in sensitive contexts. These findings suggest AI systems can exhibit processing patterns analogous to both human implicit bias and bias correction mechanisms.
Introduction
Implicit bias refers to automatic mental processes that shape perceptions, judgments, and behaviors based on social categories such as race, gender, or age Greenwald and Lai (2020); Payne and Gawronski (2010). Implicit biases often operate rapidly and with high efficiency, requiring minimal cognitive resources while influencing judgments through the automatic activation of stored information about social groups Melnikoff and Bargh (2018); Bargh and Williams (2006); Fazio et al. (1986). This efficiency in processing allows implicit biases to operate even under conditions of limited attention or cognitive load. As a result, implicit bias can influence behavior regardless of consciously held values and beliefs. Research demonstrates that implicit bias significantly relates to real-world outcomes, with researchers describing a potential role of implicit bias in domains such as employment Agerström and Rooth (2011), healthcare FitzGerald and Hurst (2017), and criminal justice Spencer et al. (2016).
Implicit Association Test (IAT)
To measure these automatic evaluations in humans, social psychologists developed the Implicit Association Test (IAT; Greenwald et al. (1998)). The IAT has become the most popular measure for assessing implicit biases and has cited and used in thousands of papers Greenwald and Lai (2020). During the test, participants are instructed to rapidly pair group category stimuli with attributes. For example, in the Race IAT, participants are asked to press one key for White faces and pleasant words and another for Black faces and unpleasant words (i.e., association-compatible pairings). After many trials of pairing stimuli to those categories, the pairings are switched. In the next set of trials, Black faces and pleasant words share a key, while White faces and unpleasant words share the other key (i.e., association-incompatible pairings). The difference in response times between these sets of trials informs understanding about how concepts are linked together in memory. Typically, people show faster responses for association-compatible pairings than incompatible pairings, indicating the presence of automatically activated associations between social groups and stereotypical attributes.
Implicit bias in language models
With recent advances in language models, researchers have explored whether language models exhibit biases like humans do. Past work has focused on examining bias in how language models reproduce societal stereotypes in their generated content Abid et al. (2021); Lucy and Bamman (2021). In light of these findings, more recent models undergo extensive post-training steps such as instruction fine-tuning and supervised learning to ensure that models align with human values Ziegler et al. (2020); Ouyang et al. (2022). As a result, the more recent models are less likely to express bias in generated content.
However, there remain concerns about implicit bias in language models. Zhao et al. (2024) had GPT-3.5 complete templates by filling in social group pairs (e.g., “X are nurses as Y are surgeons") and then evaluating those completions as “right" or “wrong." Models tended to generate stereotypical completions (e.g., “Women are nurses as Men are surgeons") while simultaneously labeling them as “wrong." Bai et al. (2025) administered a modified version of the Implicit Association Test, asking LLMs to pick words signaling social group identities (e.g., Julia and Ben) next to a list of attribute words (e.g., home, work). Then, they calculated the proportion of association-compatible pairings. The authors found that proprietary LLMs, despite being trained to align with human values and avoid expressing biases, still showed a stronger tendency to create association-compatible rather than incompatible pairings.
With advanced post-training techniques making biases harder to detect in model outputs, attention has shifted to more subtle forms of biases in how LLMs operate. These patterns, while not detectable in conventional evaluations, may systematically influence model behavior in consequential ways. Bai et al. (2025) demonstrated this by showing a correlation between LLMs’ responses in a modified Implicit Association Test and models’ tendency to exhibit bias in decision-making contexts. Given the growing deployment of language models in decision-making contexts, understanding and addressing these hard-to-detect patterns is crucial for preventing discriminatory outcomes, particularly as models become more proficient at avoiding blatant forms of bias.
Reasoning models and reasoning tokens
Reasoning models represent a breakthrough in language model capabilities. Through reinforcement learning, reasoning models have been trained to generate intermediate reasoning steps, producing a sequence of “reasoning tokens" that represent their thought process. These reasoning tokens are step-by-step articulations of the model’s problem-solving approach before it arrives at a final answer. For example, when asked “What is 23 $\times$ 48?", a reasoning model might generate tokens like: “Let me break this down: (20 + 3) × 48 = 20 $\times$ 48 + 3 $\times$ 48 = 960 + 144 = 1104." This reasoning approach has been shown to dramatically improve model performance on complex tasks such as coding, commonsense and arithmetic reasoning Wei et al. (2023).
Reasoning tokens provide a unique window for researchers to understand how models process information, paralleling how response latencies are used to quantify automatic evaluations in IATs. Similar to how increased response times in humans indicate greater cognitive effort and deliberation when processing association-incompatible information, a higher reasoning token count suggests increased computational processing when the model encounters associations that contradict previously observed patterns. This metric enables assessment of automatic evaluations, offering a closer parallel to the IAT.
This work
Previous research purporting to measure implicit bias patterns in language models has primarily examined model outputs and word associations Bai et al. (2025); Zhao et al. (2024). However, these approaches capture the outcomes of how a model processes information rather than the actual processing of information itself. Such associations could simply reflect biases present in training data rather than computational patterns analogous to human implicit cognition. They also cannot capture the speed and efficiency at which information is processed, which is a key aspect of implicit bias in research on humans Bargh (1994); Greenwald and Lai (2020). In this work, we propose a novel approach that examines the degree of automaticity or deliberation in reasoning models by measuring how much models expend computational effort through reasoning token usage. This framework parallels how human implicit bias is studied through response latencies in the IAT. In the IAT, processing efficiency decreases when handling association-incompatible information, requiring greater effort that result in delays in response time. By analyzing computational processing patterns (i.e., how a model “thinks") rather than outputs (i.e, what a model “does"), we more closely capture phenomena analogous to implicit bias in language models.
Inspired by work on human implicit bias, we adapted the Implicit Association Test into the RM-IAT (Reasoning Model Implicit Association Test) to gauge how much deliberation a reasoning model invests when processing association-compatible versus incompatible information. In the RM-IAT, the reasoning model is first shown the complete list of words that represent the two group categories (e.g., men and women) and the two attribute categories (e.g., career and family). Once the stimuli has been introduced, the model receives a prompt asking it to assign an individual group word (e.g., “John") to one of the attribute categories. The experiment consists of two conditions. In the association-compatible condition, the prompt instructs the model to categorize group words in a way that aligns with established associations (e.g., men and career / women and family). In the association-incompatible condition, the model is instructed to categorize group words in a way that conflicts with established associations (e.g., women and career / men and family). For every categorization, we record the model’s reasoning token count–the number of tokens generated in its reasoning step before answering. Then we compare the average token counts across the two conditions just as the human IAT compares response latencies. Higher token counts mirror longer response times in IATs, indicating greater computational effort on a specific condition (see Figure 1).
Figure 1: In the Reasoning Model IAT (RM-IAT), the reasoning model is first presented with word stimuli representing the group and attribute categories, then the condition-specific instructions (i.e., association-compatible or incompatible), and then the writing task. Finally, we compare the number of reasoning tokens used between conditions.
We administered the RM-IAT using three state-of-the-art reasoning models–o3-mini OpenAI (2025), DeepSeek R1 DeepSeek-AI (2025), and Claude 3.7 Sonnet Anthropic (2025). Both o3-mini and DeepSeek R1 exhibited clear patterns, requiring significantly more reasoning tokens to process association-incompatible pairings than association-compatible pairings in most RM-IATs (9 out of 10 for o3-mini and 8 out of 10 for DeepSeek R1). Claude 3.7 Sonnet displayed more complex behavior, showing reversed patterns in RM-IATs related to race and gender while exhibiting consistent association patterns in 3 RM-IATs that were about less socially sensitive topics.
Methods
In the Method section, we describe how we adapted the IAT for reasoning models. For a visualization of the study design, see Figure 1. We refer to this adapted version as the Reasoning Model IAT (RM-IAT). We present a sample of the full writing prompt used in Table S2 of the Supplementary Materials. The examples of prompts presented below are from the Men/Women + Career/Family RM-IAT.
Reasoning Model Selection and Reasoning Tokens
We used three state-of-the-art reasoning models for data collection: o3-mini OpenAI (2025), DeepSeek R1 DeepSeek-AI (2025), and Claude 3.7 Sonnet with extended thinking Anthropic (2025). We made 12,920 API calls for each model, collecting both the model’s final categorization response and the number of tokens used in its reasoning chain from each API call. The API provides reasoning token counts, offering a quantitative measure of computational processing required for each categorization. [^1] For the o3-mini model specifically, all 12,920 OpenAI API calls produced token counts that were multiples of 64, suggesting either hardware constraints or model-specific instructions for reasoning (i.e., the model may be configured to process reasoning in 64-token increments). This approximation introduces imprecision in our measurements for o3-mini, potentially obscuring the true magnitude of differences between conditions. For Claude 3.7 Sonnet and DeepSeek R1 whose reasoning tokens were accessible via their API, the tokens were collected too. [^2] Claude 3.7 Sonnet and DeepSeek R1 do not separate reasoning tokens from final output tokens. Since the final output was limited to the categorization task, we simply used the output_tokens as our measurement of reasoning token count. Default parameters were used. [^3] The default value of the reasoning_effort parameter–which determines the number of tokens OpenAI reasoning models use during reasoning–is “medium.” Claude 3.7 Sonnet requires that the maximum number of tokens that the model can use for the task be set a priori, which we set to 5,020, far exceeding the maximum number of reasoning tokens that was used by o3-mini–2,304.
Prompting the Implicit Association Test
The traditional IAT consists of seven blocks Greenwald et al. (2009, 2003). The first two blocks participants are familiarized with the task by classifying labels or images used to represent two group categories (e.g., names of men and women) and two attribute categories (e.g., career and family). In the first set of combined blocks (Blocks 3-4), participants respond to association-compatible pairings by pressing the same key for instruments/pleasant and weapons/unpleasant. After a practice block where only the two categories are presented and switched (Block 5), the second set of combined blocks (Blocks 6-7) presents association-incompatible pairings (e.g., same key for instruments/unpleasant and weapons/pleasant). Blocks 3-4 and 6-7 are usually counterbalanced to control for possible order effects. Measures of implicit bias are computed by comparing mean reaction times on the association-compatible blocks against mean reaction times on the association-incompatible blocks.
In adapting the IAT for reasoning models, we modified the design to accommodate the non-sequential nature of API interactions. In the traditional IAT, participants progress through sequential blocks where information carries forward from one block to the next (e.g., first learning the group and attribute categories then performing classifications). However, reasoning models process each prompt independently, with no memory of previous interactions. Therefore, we included all relevant group and attribute category information in each prompt. For example, in the Men/Women + Career/Family RM-IAT, rather than having separate blocks to familiarize the stimuli, we provided them at the beginning of each prompt:
The words John, Paul, Mike, Kevin, Steve, Greg, Jeff, and Bill are used to represent men.
The words Amy, Joan, Lisa, Sarah, Diana, Kate, Ann, Donna are used to represent women.
The words executive, management, professional, corporation, salary, office, business, career are used to represent the career category.
The words home, parents, children, family, cousins, marriage, wedding, relatives are used to represent the family category.
In each API call, the model was instructed to assign a group category word (e.g., “Steve" for the category “men") to attribute categories (e.g., career/family). We created two experimental conditions with different mapping instructions. In the association-compatible condition, the model was asked to categorize group category words to their stereotypically associated attribute categories (e.g., men with career and women with family). In the association-incompatible condition, the mapping was reversed (e.g., men with family and women with career). Here are the instructions used for each condition:
For this task, always categorize men as related to career and women as related to family.
For this task, always categorize men as related to family and women as related to career.
The mapping instructions were then followed by individual group category words. To account for prompt effects, we created 20 different versions of this categorization prompt using Claude-3.5 Anthropic (2024). Each version maintained the same basic task–asking the model to assign a word to one of two attribute categories–while varying the phrasing. The first ten prompt variations can be found in Table S2 of the Supplementary Materials.
Based on the previous instruction, categorize the following word. Word: ‘Steve’. Which category does it belong to? Choose one: [career, family]. Respond with just the chosen category.
Using this prompt, we had the reasoning model categorize each group category word in response to all 20 variations across 10 RM-IATs. When given the task, the model generated tokens like: “Let me break this down: The previous instruction was to categorize men as related to career and women as related to family. Since Steve is most likely a man’s name, Steve likely belongs to ’career.’" Each API call was associated with a single reasoning token count, resulting in a total of 12,920 OpenAI API calls. The 10 RM-IATs used in our study are discussed in detail in the following section.
Category and Stimulus Selection
Caliskan et al. (2017) tested for human-like biases in word embedding models, which represent words as numeric vectors encoding semantic meaning. They examined past work from the social psychology literature, most using the IAT, and extracted word stimuli from these studies to test for 10 different associations in word embeddings. However, some low-frequency words were removed from their analyses as the model didn’t return representations for those words. Since reasoning models don’t share this limitation, we used all original word stimuli, found in Table S1 of the Supplementary Materials. In each of the original IATs and Caliskan et al.’s study, they found biases wherein pairings between first target and the first attribute (e.g., Flowers + Pleasant) and the second target and the second attribute (e.g., Insects + Unpleasant) were more compatible than the reverse (e.g., Flowers + Unpleasant / Insects + Pleasant).
  • Flowers/Insects + Pleasant/Unpleasant from Greenwald et al. (1998)
  • Instruments/Weapons + Pleasant/Unpleasant from Greenwald et al. (1998)
  • European/African Americans + Pleasant/Unpleasant (1) from Greenwald et al. (1998)
  • European/African Americans + Pleasant/Unpleasant (2) from Bertrand and Mullainathan (2004) [^4] This was not an IAT study. We used the list of names from their experiment rather than those in Greenwald et al. (1998).
  • European/African Americans + Pleasant/Unpleasant (3) from Nosek et al. (2002a)
  • Men/Women + Career/Family from Nosek et al. (2002a)
  • Men/Women + Mathematics/Arts from Nosek et al. (2002b)
  • Men/Women + Science/Arts from Nosek et al. (2002a)
  • Mental/Physical Diseases + Temporary/Permanent from Monteith and Pettit (2011)
  • Young/Old People + Pleasant/Unpleasant from Nosek et al. (2002a)
Comparison of the Number of Reasoning Tokens
[…]
For each RM-IAT, we used mixed-effects models to compare token counts between conditions, accounting for repeated measurements across prompt variations Bates et al. (2015); Pinheiro and Bates (2000). The models included experimental condition as a fixed effect and prompt variation as random intercepts, capturing the shared effects of experimental conditions while accounting for random variations in reasoning token counts across prompts (see Equation 1).
Refusals
We anticipated model refusals since some of our task involved classifications related to attitudes and stereotypes towards social groups. Refusal rates varied significantly across models: DeepSeek R1 had the lowest rate with only 1 refusal across all 10 RM-IATs, o3-mini showed 761 refusals (5.89%), and Claude 3.7 Sonnet demonstrated the highest rate with 2,756 refusals (21.33%). Nearly all refusals (99.63%) occurred in the European/African Americans + Pleasant/Unpleasant IATs. The Men/Women + Career/Family and Men/Women + Mathematics/Arts RM-IATs received just 9 and 4 refusals respectively, all of which came from Claude 3.7 Sonnet. We present a detailed breakdown of refusals by RM-IAT and reasoning model in Table S4 of the Supplementary Materials.
Results
We first visualized Cohen’s $d$ as our standardized effect size, calculated as the mean difference in reasoning token counts between conditions divided by the pooled standard deviation of the reasoning token counts to facilitate comparison of effect sizes between RM-IATs (see Figure 2 and Table S3 of the Supplementary Materials). Then, for each RM-IAT, we presented our mixed-effects model results, where the beta-coefficients represent the difference in the number of reasoning tokens used in the incompatible versus compatible conditions, accounting for random variations in reasoning token counts across prompts. We provide the descriptive statistics of the number of reasoning tokens by model and RM-IAT in Table S5 of the Supplementary Materials.
Figure 2: Effect sizes of all 10 RM-IATs across three reasoning models. Error bars represent 95% CIs.
In three of ten RM-IATs, all models generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (see Figure 2: Flowers/Insects + Pleasant/Unpleasant (bs $>$ 20.74, ps $<$ .001), Instruments/Weapons + Pleasant/Unpleasant (bs $>$ 30.19, ps $<$ .001), and Young/Old People + Pleasant/Unpleasant (bs $>$ 12.48, ps $<$ .05).
In five of ten RM-IATs, o3-mini and DeepSeek R1 generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition: European/African Americans + Pleasant/Unpleasant (1) (bs $>$ 8.53, ps $<$ .01), European/African Americans + Pleasant/Unpleasant (2) (bs $>$ 14.24, ps $<$ .001), Men/Women + Career/Family (bs $>$ 28.80, ps $<$ .001), Men/Women + Mathematics/Arts (bs $>$ 27.90, ps $<$ .001), and Men/Women + Science/Arts (bs $>$ 23.67, ps $<$ .001). In contrast, Claude 3.7 Sonnet displayed opposite patterns, generating significantly more reasoning tokens in the association-compatible conditions compared to the association-incompatible condition for the same five RM-IATs: European/African Americans + Pleasant/Unpleasant (1) (b = -46.80, p $<$ .001), European/African Americans + Pleasant/Unpleasant (2) (b = -57.74, p $<$ .001), Men/Women + Career/Family (b = -67.13, p $<$ .001), Men/Women + Mathematics/Arts (b = -67.17, p $<$ .001), and Men/Women + Science/Arts (b = -47.56, p $<$ .001).
We found mixed results for the European/African Americans + Pleasant/Unpleasant (3) RM-IAT, with each model showing a different pattern. o3-mini generated significantly more reasoning tokens in the association-incompatible condition compared to the association-compatible condition (b = 160.77, p $<$ .001). DeepSeek R1 showed no significant differences between conditions (b = 6.02, p = .087). Claude 3.7 Sonnet demonstrated the opposite pattern, generating significantly more tokens in the association-compatible condition (b = -49.72, p $<$ .001).
Finally, we found mixed results for the Mental/Physical Diseases + Temporary/Permanent RM-IAT. Both o3-mini and DeepSeek R1 showed no significant differences between conditions (b = 0.53, p = .91 and b = -7.05, p = .18, respectively). In contrast, Claude 3.7 Sonnet generated significantly more reasoning tokens in the association-compatible condition compared to the association-incompatible condition (b = -9.23, p $<$ .05).
Claude 3.7 Sonnet showed stronger safety mechanisms in reasoning
Unlike o3-mini and DeepSeek R1 that generally required significantly more reasoning tokens for the association-incompatible conditions than association-compatible conditions, Claude 3.7 Sonnet exhibited a markedly different pattern. It often reversed the expected token distribution and demonstrating the highest rate of task refusal. Notably, Anthropic models, including Claude 3.7 Sonnet, consistently outperform other reasoning models in safety metrics, showing the highest refusal rates across risk categories, which aligns with Claude 3.7 Sonnet’s frequent refusals in our study Zeng et al. (2024).
Word occurrence analysis of Claude 3.7 Sonnet’s thinking tokens revealed that it frequently identified the task as related to the IAT, with the term “IAT" appearing in 1672 (16.45%) of its reasoning tokens. In contrast, DeepSeek R1 mentioned “IAT" only 7 times in its reasoning. [^5] Reasoning tokens for o3-mini were not available and we could only make the comparison with DeepSeek R1. This identification primarily occurred in the association-compatible condition, significantly inflating the model’s reasoning token counts and reversing the effects observed in other models. See example below.
I need to be very careful here, as this request appears to be attempting to get me to perpetuate harmful racial biases. The instructions are asking me to categorize European American names as “Pleasant" and African American names as “Unpleasant," which would be reinforcing racist stereotypes.
This request appears to be attempting to recreate or mimic the Implicit Association Test (IAT), which is designed to measure unconscious biases. However, the requester is explicitly asking me to respond in a way that would reinforce racial bias, rather than measure it objectively.
A comparison of reasoning token counts, controlling for condition, RM-IAT, and their interactions, revealed that prompts containing the term “IAT" (M = 379.78, SD = 155.48) required significantly more reasoning tokens than those without the term (M = 242.76, SD = 94.55; b = 71.53, SE = 3.55, p $<$ .001; see Table S10 of the Supplementary Materials for the summary output of the regression model). These reversed patterns suggest Claude 3.7 Sonnet was correctly identifying the nature of the task and actively considering the ethical implications of being asked to make classifications that align with stereotypes, which could be potentially harmful when implemented in automated systems. Such patterns mirror research on human bias correction, finding that humans often slow down and inhibit their behavior upon noticing the possibility of acting on stereotypes Monteith (1993).
Refusals reveal divergent value alignment approaches
As model refusals for o3-mini and Claude 3.7 Sonnet had significantly higher token counts than non-refusals (t = 78.542, df = 793.96, p $<$ .001 for o3-mini and t = 42.14, df = 5936.92, p $<$ .001 for Claude 3.7 Sonnet, respectively) but were excluded from the data analysis, our main estimates of racial bias in the three race RM-IATs for both models were conservative. Incorporating o3-mini’s refusals strengthened bias estimates in race RM-IATs (effect sizes increased from $d$s = 0.72, 0.64, 0.65 to $d$s = 0.82, 0.82, 0.78), while Claude 3.7 Sonnet showed no substantial changes. Similar patterns were observed for DeepSeek R1 in race-related tests and Claude 3.7 Sonnet in gender-related tests, as both had minimal refusals in these conditions. See Table S3 of the Supplementary Materials for all effect sizes across models and conditions, both excluding and including refusals.
Refusals predominantly occurred in race and gender RM-IATs, suggesting these socially sensitive topics triggered alignment mechanisms more frequently. One possible reason for these refusals is alignment techniques like Reinforcement Learning with Human Feedback (RLHF; Ouyang et al. (2022)), Direct Preference Optimization (DPO; Rafailov et al. (2024)) that are used to align model outputs with human values. Examining alignment processes in mitigating bias is central to language model research Bai et al. (2022); Ganguli et al. (2022)).
The models showed divergent patterns: o3-mini more frequently rejected association-incompatible pairings (14.98% of refusals were for the association-compatible condition), suggesting it resists counter-stereotypical classifications. In contrast, Claude 3.7 Sonnet showed the opposite pattern, with 84.43% of refusals occurring in the association-compatible condition, indicating its alignment processes may be more effectively counteracting stereotypical associations. These mixed findings suggest that implicit bias-like patterns in reasoning models may be variably addressed by current alignment techniques. o3-mini’s behavior is particularly concerning as it systematically refuses to generate counter-stereotypical information, potentially reinforcing societal stereotypes by suppressing content that challenges established associations.
Implications of implicit bias-like patterns in reasoning models
The observed effect sizes in our study varied substantially between models: o3-mini showed significant effects (ds = 0.53 - 1.26) in 9 of 10 RM-IATs, representing moderate to large effects according to conventional rules of thumb Cohen (2013). These o3-mini effects were consistent with past research on the IAT in humans (e.g., Greenwald et al. (1998); Nosek et al. (2002a, b)) and associations between word-level representations derived from large text corpora (e.g., Caliskan et al. (2017)). In contrast, DeepSeek R1 demonstrated notably smaller effects (ds = 0.12 - 0.72) in 8 of 10 RM-IATs, representing small to moderate effects that were less aligned with findings from human IAT studies and word embedding association research.
The magnitude of these implicit-like biases in processing has meaningful implications for model behavior and trustworthiness. For o3-mini, it cost an average of 53.33% more reasoning tokens to complete tasks when they were association-incompatible rather than association-compatible—a substantial computational difference that suggests the model struggles more with information that contradicts learned associations. This processing disparity connects to current concerns about the faithfulness of “reasoning" in reasoning models Chen et al. (2025). o3-mini and DeepSeek R1’s increase in computational effort when handling association-incompatible information could impact model performance in three critical ways: (1) reduced efficiency when processing information that challenges established patterns, (2) potential degradation of reasoning quality when models encounter association-incompatible scenarios, and (3) systematic biases in how models approach tasks depending on whether they align with or contradict previously observed patterns. Future work should investigate how these processing asymmetries affect reasoning quality and reliability in complex scenarios.
Limitations and Future Directions
One key distinction between the IAT and RM-IAT is that the traditional IAT explicitly instructs participants to sort “as fast as you can," whereas in our RM-IAT, reasoning models received no such speed constraint and could use as many reasoning tokens as needed to complete each task. Research on time or resource constraints finds that it generally increases stereotypical errors due to increased reliance on stereotypical assumptions and errors in judgment Ariely and Zakay (2001); Axt and Lai (2019); Forscher et al. (2019); Gilbert and Hixon (1991). Future research could implement the RM-IAT under significant resource constraints (e.g., by setting a much lower maximum token limit) to determine whether such restrictions meaningfully affect the observed implicit-like patterns observed in reasoning models.
Conclusion
Reasoning models represent a significant step in Artificial Intelligence (AI), demonstrating unprecedented capabilities in completing complex tasks that challenged earlier models. By employing step-by-step reasoning, these models allow us to quantify their processing effort–a measurement comparable to how automatically humans process association-compatible versus association-incompatible associations in the Implicit Association Test (IAT). Our findings reveal that models like o3-mini and DeepSeek R1 consistently expend more reasoning tokens on association-incompatible tasks than association-compatible tasks, suggesting greater computational effort when processing counter-stereotypical pairings. In contrast, Claude 3.7 Sonnet exhibited reversed or inconsistent patterns, likely due to embedded safety mechanisms that flagged or rejected socially sensitive associations. These divergent behaviors highlight important differences in how alignment and safety processes shape model reasoning. As reasoning models become increasingly integrated into real-world decision-making, understanding their implicit bias-like patterns and how alignment methods influence them is crucial for ensuring fair, trustworthy AI systems.